OpenStudy (anonymous):

Find the midpoint (-9, 5) (-11, -10)

OpenStudy (anonymous):

The midpoint C is located at (-10,-2.5)

OpenStudy (accessdenied):

Use the midpoint formula \[ M = \left(\frac{x_1 + x_2}{2} , \frac{y_1+y_2}{2}\right)~~~ (x_1,y_1) ~and~ (x_2,y_2) \]